OpenStudy (anonymous):

Estimate the length of one side of a square floor if the area is 820 square feet. Give the whole number that is closest to the amount.

OpenStudy (lgbasallote):

area of a square = side^2 so if the area is 820..to get the side then just take the square root what is \(\sqrt{820}?\)

OpenStudy (anonymous):

we know 30^2 = 900 so the answer must be less than 30 (it is a square floor) therefore answer is 29

OpenStudy (lgbasallote):

could be 28..that's less than 30 too